OpenStudy (anonymous):

Given f(x) = 2x-2/4 solve for f-1(3)

OpenStudy (freckles):

2x-2/4 or (2x-2)/4 ?

OpenStudy (freckles):

whatever the case is solve f(x)=3 for x and that value will be f^(-1)(3)

OpenStudy (anonymous):

So -3? @freckles

OpenStudy (freckles):

I don't think (2x-2)/4=3 gives you x=-3 could be wrong though

OpenStudy (anonymous):

my answer choices are 1 5 7 8

OpenStudy (freckles):

solve this correctly and you have your answer \[\frac{2x-2}{4}=3 \]

OpenStudy (anonymous):

thats not one of my choices

OpenStudy (freckles):

lol you have to SOLVE IT!

OpenStudy (anonymous):

I don't understand how!

OpenStudy (freckles):

undo the division by 4 by multiplying both sides by 4

OpenStudy (freckles):

\[ 4 \cdot \frac{2x-2}{4}=4 \cdot 3\]

OpenStudy (freckles):

The 4's on the left hand side will cancel since 4/4=1

OpenStudy (anonymous):

What am I putting in for x?

OpenStudy (freckles):

no no you are solving for x

OpenStudy (freckles):

\[\cancel{4} \cdot \frac{2x-2}{\cancel{4}}=4 \cdot 3 \]

OpenStudy (freckles):

Come on I know you can solve 2x-2=12

OpenStudy (freckles):

give it a try

OpenStudy (anonymous):

2x=14

OpenStudy (freckles):

ok one last step

OpenStudy (anonymous):

7 right?

OpenStudy (freckles):

\[f(7)=3 \text{ implies } f^{-1}(3)=7\]

OpenStudy (anonymous):

Thank you! Can you help me with two more?

OpenStudy (anonymous):

@freckles